# Kernel Three Pass Regression Filter

### v3

@@ Abstract
We forecast a single time series using a high-dimensional set of predictors. When predictors share common underlying dynamics, a latent factor model estimated by the Principal Component method effectively characterizes their co-movements. These latent factors succinctly summarize the data and aid in prediction, mitigating the curse of dimensionality. However, two significant drawbacks arise: (1) not all factors may be relevant, and utilizing all of them in constructing forecasts leads to inefficiency, and (2) typical models assume a linear dependence of the target on the set of predictors, which limits accuracy. We address these issues through a novel method: Kernel Three-Pass Regression Filter. This method extends a supervised forecasting technique, the Three-Pass Regression Filter, to exclude irrelevant information and operate within an enhanced framework capable of handling nonlinear dependencies. Our computationally efficient method demonstrates strong empirical performance, particularly over longer forecast horizons.
@@ 1 Introduction
In recent years, the surge in high-dimensional datasets across fields like economics has ushered in new opportunities and challenges. A paramount issue is the ‘curse of dimensionality,’ which undermines the effectiveness of traditional finite-dimensional estimation methods. Most modeling techniques applied to high-dimensional data assume the existence of a low-dimensional structure that effectively summarizes the data. One stylized feature of high-dimensional economic datasets is the presence of high and pervasive collinearity among variables, leading researchers to posit a data-generating process that assumes all variables are a function of a few latent factors. This formulation is commonly referred to as the factor model. A vast amount of literature focuses on using this latent factor structure for forecasting applications. A typical example is found in diffusion index models (Stock & Watson, (2002b)), where latent factors are derived from a high-dimensional set of variables using Principal Components (hereafter, PC) method. These factors are subsequently utilized to forecast a target variable. A limitation of this PC-based factor estimation is its unsupervised nature, i.e., no information from the target variable is incorporated.
@@
Since the primary goal is to forecast a target rather than estimate the underlying factor structure, introducing a degree of supervision can be beneficial. This can help filter out irrelevant information from the predictor set, thus enhancing the predictive accuracy. This can be done in different ways: using soft and hard thresholding methods to remove predictors with no predictive content, as in Bai & Ng, (2008), or assigning varying weights to predictors based on their predictive capabilities for the target (see, for example, Huang et al., (2022)), or estimate the subset of factors that exhibit predictive power for the target rather than the complete set of factors that drive the predictors, as in Kelly & Pruitt, (2015).
@@
The aforementioned models, whether utilizing PCA or supervised factor models, are predicated on the convenient assumption of linearity. However, as underscored in Goulet Coulombe et al., (2022), non-linearity often characterizes many predictive relationships in economics, particularly over extended time horizons and within data-rich environments.
@@
Various approaches have been proposed to integrate non-linearity into factor models. For instance, squared principal components (PCs) or principal component squared ($PC^{2}$) as seen in Bai & Ng, (2008), sufficient forecasting by Fan et al., (2017), the kernel trick to estimate factors (Kutateladze, (2022)) among others. However, these approaches have limited supervision in the prediction process, if any. For example, Fan et al., (2017) estimates factors through an unsupervised method (PC) and then derives sufficient indices using these PCs. Similarly, Kutateladze, (2022) essentially applies kernel PCA (an unsupervised method) to estimate the set of factors driving a higher-dimensional space obtained by lifting the set of predictors through the kernel method. In Bai & Ng, (2008), a very particular form of non-linearity (quadratic) is examined, which is somewhat ad hoc. Although they employ thresholding methods to reduce predictors to a smaller set, their screening method, however, may still encounter challenges in filtering relevant factors within this subset, leading to inefficient forecasts.
Our paper incorporates both non-linearity and supervision by introducing a novel kernel three-pass regression filter. Our approach essentially applies the three-pass filter (hereafter 3PRF) proposed by Kelly & Pruitt, (2015) to a transformed set of predictors. We adopt the lifting concept similar to Kutateladze, (2022), but instead of employing an unsupervised method like kernel PCA, we utilize a supervised method to estimate factors relevant to the target variable.
The table below summarizes our discussion by listing some popular methods [^1] The entries in this table are some of the most popular forecasting methods used in econometric literature. However, by no means do they form an exhaustive set. in literature and how this paper is placed among them
Linear Non-Linear
Unsupervised PC kernel PCA, Sq-PC, $PC-sq$
Supervised 3PRF This Paper
Table 1: Factor Model Based Forecasting Methods
@@
The paper proceeds as follows. Section 2 provides a brief introduction to Kernel methods. Section 3 introduces our estimator and discusses its similarity with the estimator of Kelly & Pruitt, (2015). We also list a set of assumptions that ensure the theoretical properties of our estimators, which are given in the subsequent section 4. We present our empirical results in sections 5 and 6 and conclude in section 7. Mathematical proofs and implementation details are given in the appendix.
@@
We use $\boldsymbol{y}$ to denote the $\text{T}\times 1$ vector of the target variable, i.e. $\boldsymbol{y}=(y_{h},y_{h+1}\dots y_{t+h})$. We have $N$ predictors with $T$ observations for each predictor. The cross section of predictors at a time $t$ is given by the $N\times 1$ vector $\boldsymbol{x}_{t}$. Similarly, the vector of temporal observations of a predictor $i$ is given by $\mathbf{x}_{i}$. We stack the predictors in a ${T}\times{N}$ matrix $\boldsymbol{X}$, $\boldsymbol{X}=$ $\left(\boldsymbol{x}_{1}^{\prime},\boldsymbol{x}_{2}^{\prime},\ldots,\boldsymbol{x}_{T}^{\prime}\right)^{\prime}=\left(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\right)$. We have $L$ proxies which we stack in a ${T}\times\ {L}$ matrix $\boldsymbol{Z}=\left(\boldsymbol{z}_{1}^{\prime},\boldsymbol{z}_{2}^{\prime},\ldots,\boldsymbol{z}_{T}^{\prime}\right)^{\prime}$. The demeaning matrix $\boldsymbol{J}_{T}\equiv\boldsymbol{I}_{T}-\frac{1}{T}\mathbf{\iota}_{T}\mathbf{\iota}_{T}^{\prime}$, where $\boldsymbol{I}_{T}$ is the $T$-dimensional identity matrix and $\mathbf{\iota}_{T}$ the $T$-vector of ones. For matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ of conformable dimensions, $\boldsymbol{S}_{UV}\equiv\boldsymbol{U}^{\prime}\boldsymbol{J}_{T}\boldsymbol{V}$. For the transformed set of predictors $\varphi(\boldsymbol{X})$, $\varphi_{j}(\mathbf{x})$ denotes the $j^{th}$ observation. $\varphi(\boldsymbol{X})=$ $\left(\varphi_{1}(\mathbf{x}),\varphi_{2}(\mathbf{x}),\ldots,\varphi_{M}(\mathbf{x})\right)=$ $\left(\varphi(\boldsymbol{x}_{1})^{\prime},\varphi(\boldsymbol{x}_{2})^{\prime},\ldots,\varphi(\boldsymbol{x}_{T})^{\prime}\right)^{\prime}$. Stochastic orders are denoted by the usual ${O}_{p}$ and ${o}_{p}$. For a matrix, $\boldsymbol{O}_{p}$ and $\boldsymbol{o}_{p}$ denotes the element wise stochastic order, i.e., a matrix is said to be $\boldsymbol{O}_{p}(1)$ or $\boldsymbol{o}_{p}(1)$ if all it’s elements are ${O}_{p}(1)$ or ${o}_{p}(1)$ respectively.
@@ 2 Kernel Method
Let $\varphi:\boldsymbol{X}\rightarrow\mathcal{F}$ denote a transformation of the original data into a higher-dimensional space [^2] Precisely, we are referring to a Hilbert space where the inner product of the vectors is well-defined. containing the original set of predictors and their non-linear transformations. Methods such as principal components or the three-pass regression filter depend on the input $\boldsymbol{X}$ only through the $T\times T$ matrix of dot products $\boldsymbol{X}^{\prime}\boldsymbol{X}$. Applying these methods to the transformed predictors $\varphi(\boldsymbol{X})$ would therefore require computing the inner product $\varphi(\boldsymbol{X})^{\prime}\varphi(\boldsymbol{X})$. This computation can be cumbersome or infeasible [^3] When the transformed space is infinite-dimensional. Here, the kernel trick proves to be handy, allowing us to calculate inner products within the transformed space without requiring explicit knowledge of $\varphi$. A valid [^4] A positive semi-definite kernel as discussed in Appendix-A.2. kernel corresponds to an inner product of features $\varphi(\boldsymbol{X})$, where the analytical form of the function $\varphi(\cdot)$ may be unknown, but it is guaranteed to exist by Mercer’s theorem (Appendix-A.2). Hence, using a Kernel function to compute inner products within a method is akin to performing the estimation exercise(implicitly) on the set of transformed features. In the Supplementary appendix-B.1, we illustrate how different kernel functions correspond to the inner products of transformed inputs.
Utilizing a transformed set of predictors provides a significant advantage, as many nonlinear relationships can be reformulated as linear relationships in the appropriately transformed space. As an illustration, consider the following example. We generate two variables $X$ and $Y$ from uniform distribution $U[-2,2]$. Define a binary variable $z$ as:
$z=\begin{cases}1&\text{if }X^{2}+Y^{2}\leq 2\\ -1&\text{otherwise}\end{cases}$
As shown in the figure-1 (left), a linear boundary cannot separate the two classes of variable $z$. However, upon transforming the original spaces $X$ and $Y$ to $\varphi_{1}(X)=\sin^{2}(X)$ and $\varphi_{2}(Y)=\cos^{2}(Y)$ respectively, we find that the two classes can be easily distinguished as seen in figure-1(right). The blue points are in class 1, and the red ones are in class -1.
@@
For the sake of simplicity, this example illustrates the transformation of a two-dimensional input $\boldsymbol{W}=(X,Y)$ into a two-dimensional feature space. The transformed space $\varphi(\boldsymbol{W})$ is typically high-dimensional and potentially infinite-dimensional. Transformation to a higher-dimensional space makes a large set of non-linear forms available, rendering the discovery of a nonlinear relationship very likely.
@@ 3 The Estimator
We delineate the three regression passes that we use to construct our forecast. The first two passes, as explained below, are not feasible in practice, whilst the eventual closed-form solutions are. Nonetheless, these steps offer valuable insights into the underlying process of our estimator and elucidate its similarity to the well-known linear three-pass filter proposed by Kelly & Pruitt, (2015).
@@
$\varphi$ maps our N-dimensional predictors $\boldsymbol{x}_{t}$ to a higher M-dimensional space. Assumption 1 endows this transformed set of predictors with a factor structure. An underlying factor structure among $\boldsymbol{X}$ implies the existence of a low dimensional plane, projection onto which explains maximal variation in the predictors. An equivalent interpretation of a linear factor structure on $\varphi(\boldsymbol{x}_{t})$ would be the existence of a lower dimensional manifold which explains maximum variation in $\boldsymbol{x}_{t}$. This manifold’s basis comprises a few uni-dimensional orthogonal projections of $\varphi(\boldsymbol{x}_{t})$.
The infeasible three-passes are summarized in 2 below.

### v4

@@ Abstract
We forecast a single time series using a high-dimensional set of predictors. When predictors share common underlying dynamics, a latent factor model estimated by the Principal Component method effectively characterizes their co-movements. These latent factors succinctly summarize the data and aid in prediction, mitigating the curse of dimensionality. However, two significant drawbacks arise: (1) not all factors may be relevant, and utilizing all of them in constructing forecasts leads to inefficiency, and (2) typical models assume a linear dependence of the target on the set of predictors, which limits accuracy. We address these issues through a novel method: Kernel Three-Pass Regression Filter. This method extends a supervised forecasting technique, the Three-Pass Regression Filter, to exclude irrelevant information and operate within an enhanced framework capable of handling nonlinear dependencies. Our method is computationally efficient and demonstrates strong empirical performance, particularly over longer forecast horizons.  [^1] We are grateful to the Editor, the three anonymous referees, and the participants at the 2024 California Econometrics Conference, the 2025 World Congress of the Econometric Society in Seoul, the 34th Midwest Econometrics Group Conference in Kentucky (USA), the 2024 European Winter Meeting of the Econometric Society in Palma de Mallorca (Spain), and the seminar at the University of California, Riverside, for their valuable comments and suggestions.We also thank the Econometric Society for selecting our paper among the top 10% of contributions presented at the 2025 World Congress for the USD 2,000 fellowship award.
@@ 1 Introduction
In recent years, the surge in high-dimensional datasets across fields like economics has ushered in new opportunities and challenges. A paramount issue is the ‘curse of dimensionality,’ which undermines the effectiveness of traditional finite-dimensional estimation methods. Most modeling techniques applied to high-dimensional data assume the existence of a low-dimensional structure that effectively summarizes the data. One stylized feature of high-dimensional economic datasets is the presence of high and pervasive collinearity among variables, leading researchers to posit a data-generating process that assumes all variables are a function of a few latent factors. This formulation is commonly referred to as the factor model. A vast amount of literature focuses on using this latent factor structure for forecasting applications. A typical example is found in diffusion index models (stock2002diff), where latent factors are derived from a high-dimensional set of variables using Principal Components (hereafter, PC) method. These factors are subsequently utilized to forecast a target variable. A limitation of this PC-based factor estimation is its unsupervised nature, i.e., no information from the target variable is incorporated.
@@
Since the primary goal is to forecast a target rather than estimate the underlying factor structure, introducing a degree of supervision can be beneficial. This can help filter out irrelevant information from the predictor set, thus enhancing the predictive accuracy. This can be done in different ways: using soft and hard thresholding methods to remove predictors with no predictive content, as in bai2008, or assigning varying weights to predictors based on their predictive capabilities for the target (see, for example, scaledpca), or estimate the subset of factors that exhibit predictive power for the target rather than the complete set of factors that drive the predictors, as in kelly.
@@
The aforementioned models, though offering a supervised alternative to the unsupervised nature of Principal Components, nevertheless rest on the convenient assumption of linearity. However, as underscored in ml_jae, non-linearity often characterizes many predictive relationships in economics, particularly over extended time horizons and within data-rich environments.
@@
Various approaches have been proposed to integrate non-linearity into factor models. For instance, squared principal components (PCs) or principal component squared ($PC^{2}$) as seen in bai2008, sufficient forecasting by fanSDR, the kernel trick to estimate factors (Varlam) among others. However, these approaches have limited supervision in the prediction process, if any. For example, fanSDR estimates factors through an unsupervised method (PC) and then derives sufficient indices using these PCs. Similarly, Varlam essentially applies kernel PCA (an unsupervised method) to estimate the set of factors driving a higher-dimensional space obtained by lifting the set of predictors through the kernel method. In bai2008, a very particular form of non-linearity (quadratic) is examined, which is somewhat ad hoc. Although they employ thresholding methods to reduce predictors to a smaller set, their screening method, however, may still encounter challenges in filtering relevant factors within this subset, leading to inefficient forecasts.
Our paper incorporates both non-linearity and supervision by introducing a novel kernel three-pass regression filter. The approach applies the three-pass regression filter of kelly to a non-linear transformation of the predictor set. While we adopt the lifting idea used in Varlam, we depart from unsupervised techniques such as kernel PCA and instead employ a supervised procedure to estimate factors obtained from non-linear transformations of the predictors that are relevant for the target variable. Our results align with the findings in ml_jae; over long horizons, where non-linear predictive relations become more pronounced, methods that capture non-linearity, such as ours, consistently outperform their linear counterparts. However, other non-linear approaches do not perform as well. Although they display the expected pattern of improving relative to linear models at longer horizons, the gains are substantially smaller. We attribute the superior performance of our approach to the combination of non-linearity and supervision. Importantly, the non-linear gains at longer horizons do not come at the cost of short-horizon underperformance; while it may not always be the best at very short horizons, it remains close to the best and in several cases even surpasses competing methods.
@@
The paper proceeds as follows. Section 2 provides a brief introduction to Kernel methods. Section 3 introduces our estimator and discusses its similarity with the estimator of kelly. We also list a set of assumptions that ensure the theoretical properties of our estimators, which are given in the subsequent section 4. We present our empirical results in sections 5 and 6 and conclude in section 7. Mathematical proofs and implementation details are given in the appendix.
@@
We use $\boldsymbol{y}$ to denote the $T\times 1$ vector of the target variable, i.e.
$\boldsymbol{y}=(y_{h},y_{h+1},\dots,y_{T+h})^{\prime}.$
We have $N$ predictors with $T$ observations for each predictor. The cross section of predictors at time $t$ is given by the $N\times 1$ vector $\boldsymbol{x}_{t}$, while the temporal observations of predictor $i$ are given by the vector $\mathbf{x}_{i}\in\mathbb{R}^{T}$. Stacking the predictors yields the $T\times N$ matrix
$\boldsymbol{X}=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T})^{\prime}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}),$
so that the $t$-th row of $\boldsymbol{X}$ is the cross section $\boldsymbol{x}_{t}^{\prime}\in\mathbb{R}^{1\times N}$, and the $i$-th column contains the temporal observations $\mathbf{x}_{i}\in\mathbb{R}^{T\times 1}$ of predictor $i$.
We also have $L<\infty$ proxies, collected in the $T\times L$ matrix
$\boldsymbol{Z}=(\boldsymbol{z}_{1},\boldsymbol{z}_{2},\ldots,\boldsymbol{z}_{T})^{\prime},$
where each $\boldsymbol{z}_{t}\in\mathbb{R}^{L}$ is the vector of all $L$ proxies observed at time $t$.
We use $\varphi$ to denote a transformation $\varphi:\mathbb{R}^{N}\to\mathbb{R}^{M}$, which transforms each $\boldsymbol{x}_{t}$, at time $t$, into an $M\times 1$ vector $\varphi(\boldsymbol{x}_{t})$. Applying $\varphi$ to the vector of predictors at each time point and stacking the resulting vectors yields the transformed predictor matrix,
$\varphi(\boldsymbol{X})=\big(\varphi(\boldsymbol{x}_{1}),\varphi(\boldsymbol{x}_{2}),\dots,\varphi(\boldsymbol{x}_{T})\big)^{\prime}=\big(\varphi_{1}(\mathbf{x}),\varphi_{2}(\mathbf{x}),\dots,\varphi_{M}(\mathbf{x})\big)\in\mathbb{R}^{T\times M},$
where the $t$-th row of $\varphi(\boldsymbol{X})$ is the transformed cross section $\varphi(\boldsymbol{x}_{t})^{\prime}\in\mathbb{R}^{1\times M}$ at time $t$, and the $j$-th column $\varphi_{j}(\mathbf{x})\in\mathbb{R}^{T\times 1}$ represents the vector of temporal values of the $j$-th transformed feature across times $\{1,\dots,T\}$. $\boldsymbol{I}_{L}$ denotes an identity matrix of dimesnion $L$. We use $\boldsymbol{J}_{T}$ to denote the demeaning matrix, i.e., $\boldsymbol{J}_{T}\equiv\boldsymbol{I}_{T}-\frac{1}{T}\mathbf{\iota}_{T}\mathbf{\iota}_{T}^{\prime}$, and for conformable matrices $\boldsymbol{U}$ and $\boldsymbol{V}$ we define $\boldsymbol{S}_{UV}\equiv\boldsymbol{U}^{\prime}\boldsymbol{J}_{T}\boldsymbol{V}.$. Stochastic orders are denoted by the usual ${O}_{p}$ and ${o}_{p}$. For a matrix, $\boldsymbol{O}_{p}$ and $\boldsymbol{o}_{p}$ denotes the element wise stochastic order, i.e., a matrix is said to be $\boldsymbol{O}_{p}(1)$ or $\boldsymbol{o}_{p}(1)$ if all it’s elements are ${O}_{p}(1)$ or ${o}_{p}(1)$ respectively.
@@ 2 Kernel Method
We consider a transformation $\varphi:\mathbb{R}^{N}\to\mathbb{R}^{M}$ that maps the original set of predictors $\boldsymbol{x}_{t}$ into a vector $\varphi(\boldsymbol{x}_{t})$ in a (typically higher- or potentially infinite-) dimensional space. These transformed vectors are often called features. Stacking the transformed predictors across all times yields the matrix
$\varphi(\boldsymbol{X})=\big(\varphi(\boldsymbol{x}_{1}),\varphi(\boldsymbol{x}_{2}),\dots,\varphi(\boldsymbol{x}_{T})\big)^{\prime}\in\mathbb{R}^{T\times M}.$
Methods such as principal components and 3PRF, depend on the predictors only through the $T\times T$ matrix of dot products $\boldsymbol{X}\boldsymbol{X}^{\prime}$. Extending these methods to the transformed set of predictors would require computing the Gram matrix $\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}$, which can be computationally expensive when $M$ is large or infeasible if the transformed features lie in an infinite-dimensional space.
To address this, we introduce a kernel function $\mathcal{K}(\boldsymbol{x}_{t},\boldsymbol{x}_{s})$, which computes the inner product of the transformed predictors indirectly as
$\mathcal{K}(\boldsymbol{x}_{t},\boldsymbol{x}_{s})=\varphi(\boldsymbol{x}_{t})^{\prime}\varphi(\boldsymbol{x}_{s}).$ (2.1)
This approach, known as the ‘kernel trick’, allows us to compute the inner products in the feature space without explicitly evaluating the transformation $\varphi$. Stacking these evaluations across all times yields the kernel (Gram) matrix
$\mathcal{K}(\boldsymbol{X},\boldsymbol{X}^{\prime})=\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\in\mathbb{R}^{T\times T},$
whose $(t,s)$-th entry corresponds to $\mathcal{K}(\boldsymbol{x}_{t},\boldsymbol{x}_{s})$ as defined in (2.1). A function $\mathcal{K}$ is called a valid kernel if, for any collection of inputs $\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T}\}$, the resulting matrix $\mathcal{K}(\boldsymbol{X},\boldsymbol{X}^{\prime})$ is symmetric and positive semi-definite. By Mercer’s theorem (see Appendix A.2), every such valid kernel is associated with a Hilbert space $\mathcal{H}$ and a mapping $\varphi:\mathbb{R}^{N}\to\mathcal{H}$ such that the equality in (2.1) holds. In other words, a valid kernel can always be interpreted as the inner product between transformed observations in some feature space. [^2] In the Online Appendix-B.1, we illustrate how different kernel functions correspond to the inner products of transformed inputs.
@@
Working in a transformed space provides a substantial advantage: many nonlinear relationships among the original predictors can become linear in the appropriately transformed set of predictors. As a stylized example, consider the illustration in Figure 1. We generate two variables $X$ and $Y$ from the uniform distribution U[2,2]U[-2,2] and define a binary variable $z$ as
$z=\begin{cases}1&\text{if }X^{2}+Y^{2}\leq 2,\\ -1&\text{otherwise.}\end{cases}$
In the original predictor space (Figure 1, left), the two classes of $z$ cannot be separated by a linear boundary. However, after transforming the inputs to $\varphi_{1}(X)=\sin^{2}(X)$ and $\varphi_{2}(Y)=\cos^{2}(Y)$, the classes become linearly separable in the transformed feature space (Figure 1, right), where blue and red points represent the two classes. Although this example involves only a two-dimensional input $\boldsymbol{W}=(X,Y)$ transformed into a two-dimensional feature space, in practice $\varphi(\boldsymbol{X})$ can be high- or even infinite-dimensional. Such transformations typically introduce a large set of nonlinear basis functions of the original predictors, making it more likely that an underlying nonlinear relationship in the data becomes linearly representable in the transformed space.
@@ 3 The Estimator
We delineate the three regression passes that we use to construct our forecast. The first two passes, as explained below, are not feasible in practice, whilst the eventual closed-form solutions are. Nonetheless, these steps offer valuable insights into the underlying process of our estimator and elucidate its similarity to the well-known linear three-pass filter proposed by kelly.
@@
The mapping $\varphi$ transforms the original $N$-dimensional predictors $\boldsymbol{x}_{t}$ into an $M$-dimensional set of features, where $M$ depends on the input dimension $N$. In general, for any (fixed) valid kernel, $M$ increases with $N$, meaning that adding more predictors weakly increases the dimension of the (implied) transformed feature space for a given Kernel. For example, under a second-degree polynomial kernel, the transformed feature vector includes all original linear terms, all squared terms, and all pairwise products of the predictors, so that the total number of features is $M=N+N(N+1)/2$, which grows with NN. [^3] See the Online Appendix for more details.
Assumption 1 endows the transformed set of predictors with a factor structure. The coefficient vector $\boldsymbol{\beta}_{f}$ corresponds to the relevant factors $\boldsymbol{f}_{t}$ in the main predictive equation for $y_{t+h}$; being nonzero, it indicates that these factors directly influence the outcome, whereas the coefficients on the irrelevant factors $\boldsymbol{g}_{t}$ are zero. The dimensions $K_{f}$, $K_{g}$, and $L$ are all fixed and remain small relative to $N$.
The factor loading matrix for the transformed predictors $\boldsymbol{\Phi}$ has dimension $M\times K$, mapping the $K$ latent factors to the $M$ transformed features. Individual factor loadings $\boldsymbol{\phi}_{k}\in\mathbb{R}^{M}$ correspond to the weights of factor $k$ across the $M$ transformed features, with $\boldsymbol{\Phi}=[\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{K}]$. The submatrix $\boldsymbol{\Phi}_{f}$ is $M\times K_{f}$ and contains the loadings on the relevant factors, while $\boldsymbol{\Phi}_{g}$ is $M\times K_{g}$ and contains the loadings on the irrelevant factors. Similarly, the loading matrix for the proxies, $\boldsymbol{\Lambda}$, has dimension $L\times K$, mapping the $K$ latent factors to the $L$ proxy variables. Its submatrices $\boldsymbol{\Lambda}_{f}$ and $\boldsymbol{\Lambda}_{g}$ have dimensions $L\times K_{f}$ and $L\times K_{g}$, respectively, and are defined analogously, capturing the contribution of the relevant and irrelevant factors to each proxy.
The proxy variables $\boldsymbol{z}_{t}$ are generated by the underlying factors together with idiosyncratic proxy noise. Overall, the framework parallels kelly, with the key distinction that the factor structure is applied to transformed (and potentially nonlinear) features rather than the raw predictor space.
Our estimator is characterized by three infeasible three-passes, which are summarized in Table 1.